There were 423 people riding the train. 273 more people got on the train, and 93 got off at the next station. At the third station half of all the people got off the train. How many people are left on the train?
Answer by skippo(9)

There were 423 people riding the train. 273 more people got on the train, and 93 got off at the next station. At the third station half of all the people got off the train. How many people are left on the train?
solution:
423 people were present
1st station: 273 more people got on, so we have 423+273 =696
2nd station: 93 people got off, so we have 696-93 =603
3rd station: half of the people got off, so we have 1/2x603 = 301.5
therefore, approximately 301 people are left in the train.